• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name jroche

    Log entry time 03:41:17 on October15,2010

    Entry number 334960

    keyword=LHRS set for 2.607 GeV

    previous setting was 2.4 Gev, so we cycled Q2 and Q3.
    Had to lock the NMR on probe B.

    ready to go.

    FIGURE 1

    FIGURE 2